# TimeMixer++: A General Time Series Pattern Machine for Universal Predictive Analysis

### v1

@@ 1 Introduction
At the core of TSPMs is their ability to recognize and generalize time series patterns inherent in time series data, enabling the model to uncover meaningful temporal structures and adapt to varying time series task scenarios. A line of research (Lai et al., 2018b; Zhao et al., 2017) has utilized recurrent neural networks (RNNs) to capture sequential patterns. However, these methods often struggle to capture long-term dependencies due to limitations like Markovian assumptions and inefficiencies. Temporal convolutional networks (TCNs) (Wu et al., 2023; Wang et al., 2023; Liu et al., 2022a) efficiently capture local patterns but face challenges with long-range dependencies (e.g., seasonality and trends) because of their fixed receptive fields. While some approaches reshape time series into 2D tensors based on frequency domain information (Wu et al., 2023) or downsample the time domain (Liu et al., 2022a), they fall short in comprehensively capturing long-range patterns. In contrast, transformer-based architectures (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2022b) leverage token-wise self-attention to model long-range dependencies by allowing each token to attend to all others, overcoming the limitations of fixed receptive fields. However, unlike language tasks where tokens usually belong to distinct contexts, time series data often involve overlapping contexts at a single time point, such as daily, weekly, and seasonal patterns occurring simultaneously. This overlap makes it difficult to represent time series patterns effectively as tokens, posing challenges for transformer-based models in fully capturing the relevant temporal structures.
@@ 2 Related Work
Time Series Analysis. A pivotal aspect of time series analysis is the ability to extract diverse patterns from various time series while building powerful representations. This challenge has been explored across various model architectures. Traditional models like ARIMA (Anderson & Kendall, 1976) and STL (Cleveland et al., 1990) are effective for periodic and trend patterns but struggle with non-linear dynamics. Deep learning models, such as those by (Lai et al., 2018b) and (Zhao et al., 2017), capture sequential dependencies but face limitations with long-term dependencies. TCNs (Franceschi et al., 2019) improve local pattern extraction but are limited in capturing long-range dependencies. TimesNet (Wu et al., 2023) enhances long-range pattern extraction by treating time series as 2D signals, while MLP-based methods (Zeng et al., 2023; Ekambaram et al., 2023) offer simplicity and effectiveness. Transformer-based models like PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024) leverage self-attention to model long-range dependencies, demonstrating good forecasting performance. Given the strengths and limitations discussed above, there is a growing need for a TSPM capable of effectively extracting diverse patterns, adapting to various time series analytical tasks, and possessing strong generalization capabilities. As illustrated in Figure 1, TimeMixer++ meets this requirement by constructing robust representational capabilities, thereby demonstrating its potential for universal time series analysis.
@@
Hierarchical Time Series Modeling. Numerous methodologies have been advanced utilizing specialized deep learning architectures for time series analysis, with an emphasis on the decomposition and integration of temporal patterns. For example, several studies (Wu et al., 2021; Wang et al., 2024; Zhou et al., 2022b; Wang et al., 2024) utilize moving averages to discern seasonal and trend components, which are subsequently modeled using attention mechanisms (Wu et al., 2021; Zhou et al., 2022b), convolutional networks (Wang et al., 2023), or hierarchical MLP layers (Wang et al., 2024). These components are individually processed prior to aggregation to yield the final output. Nonetheless, such approaches frequently depend on predefined and rigid operations for the disentanglement of seasonality and trends, thereby constraining their adaptability to complex and dynamic patterns. In contrast, as depicted in Figure 2, we propose a more flexible methodology that disentangles seasonality and trend directly within the latent space via dual-axis attention, thereby enhancing adaptability to a diverse range of time series patterns and task scenarios. Furthermore, by adopting a multi-scale, multi-resolution analytical framework (Mozer, 1991; Harti, 1993), we facilitate hierarchical interaction and integration across different scales and resolutions, substantially enhancing the effectiveness of time series modeling.
@@ 3.2 MixerBlock
where $\mathbf{A}=\{A_{f_{1}},\cdots,A_{f_{K}}\}$ represents the unnormalized amplitudes, $\{f_{1},\cdots,f_{K}\}$ are the top-$K$ frequencies, and $p_{k}=\left\lceil\frac{T}{f_{k}}\right\rceil,k\in\{1,\dots,K\}$ denotes the corresponding period lengths. Each 1D time series $\mathbf{x}_{m}^{l}$ is then reshaped into $K$ 2D images as follows:
@@
Time Image Decomposition. Time series patterns are inherently nested, with overlapping scales and periods. For example, weekly sales data reflects both daily shopping habits and broader seasonal trends. Conventional methods (Wu et al., 2021; Wang et al., 2024) use moving averages across the entire series, often blurring distinct patterns. To address this, we utilize multi-resolution time images, where each image $\mathbf{z}_{m}^{(l,k)}\in\mathbb{R}^{p_{k}\times f_{\text{m,k}}\times d_{\text{model}}}$ encodes a specific scale and period, enabling finer disentanglement of seasonality and trend. By applying 2D convolution to these images, we capture long-range patterns and enhance temporal dependency extraction. Columns in each image correspond to time series segments within a period, while rows represent consistent time points across periods, facilitating dual-axis attention: column-axis attention ($\operatorname{Attention_{col}}$) captures seasonality within periods, and row-axis attention ($\operatorname{Attention_{row}}$) extracts trend across periods. Each axis-specific attention focuses on one axis, preserving efficiency by transposing the non-target axis to the batch dimension. For column-axis attention, queries, keys, and values $\mathbf{Q}_{\text{col}},\mathbf{K}_{\text{col}},\mathbf{V}_{\text{col}}\in\mathbb{R}^{f_{\text{m,k}}\times d_{\text{model}}}$ are computed via 2D convolution, which are shared across all images, and similarly for row-axis attention $\mathbf{Q}_{\text{row}},\mathbf{K}_{\text{row}},\mathbf{V}_{\text{row}}$. The seasonal and trend components are then computed as:
@@
Multi-scale Mixing. For each period $p_{k}$, we obtain $M+1$ seasonal time images and $M+1$ trend time images, denoted as $\{\mathbf{s}_{m}^{(l,k)}\}_{m=0}^{M}$ and $\{\mathbf{t}_{m}^{(l,k)}\}_{m=0}^{M}$, respectively. The 2D structure allows us to model both seasonal and trend patterns using 2D convolutional layers, which are more efficient and effective at capturing long-term dependencies than traditional linear layers (Wang et al., 2024). For multi-scale seasonal time images, longer patterns can be interpreted as compositions of shorter ones (e.g., a yearly rainfall pattern formed by monthly changes). Therefore, we mix the seasonal patterns from fine-scale to coarse-scale. To facilitate this bottom-up information flow, we apply the 2D convolutional layers at the $m$-th scale in a residual manner, formalized as:
@@ 4.1.1 Long-term forecasting
Setups. Long-term forecasting is pivotal for strategic planning in areas such as weather prediction, traffic management, and energy utilization. To comprehensively assess our model’s effectiveness over extended periods, we perform experiments on 8 widely-used real-world datasets, including the four subsets of the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), as well as Weather, Solar-Energy, Electricity, and Traffic, consistent with prior benchmarks set by Zhou et al. (2021b); Wu et al. (2021); Liu et al. (2022a)
Table 1: Long-term forecasting results. We average the results across 4 prediction lengths: $\{96,192,336,720\}$. For consistency, we maintain a fixed input length of $96$ throughout all the experiments. Avg means further averaged by subsets. The best performance is highlighted in red, and the second-best is underlined. Full results can be found in Appendix H.
@@ 4.1.2 Univariate short-term forecasting
Table 2: Univariate short-term forecasting results, averaged across all M4 subsets. Prediction lengths are in $[6,48]$, with input lengths set to twice the prediction length. Lower SMAPE, MASE, and OWA values indicate better performance. Auto. and FED. indicates Autoformer and FEDformer, while Stationary denotes the Non-stationary Transformer. Full results are available in Appendix H
@@
Results. Table 2 demonstrates that TimeMixer++ significantly outperforms state-of-the-art models across all metrics. Compared to iTransformer, it reduces SMAPE by 9.7% and MASE by 15.7%, with even larger improvements over TiDE, achieving up to a 23.3% reduction in MASE. Additionally, TimeMixer++ records the lowest OWA, outperforming TimesNet by 3.5% and iTransformer by 11.6%, setting a new standard for short-term forecasting.
@@ 4.1.3 Multivariate short-term forecasting
Table 3: Results of multivariate short-term forecasting, averaged across all PEMS datasets. The input sequence length is set to 96, while the prediction length is 12. Lower values for MAE, MAPE, and RMSE indicate better performance. Full results can be found in Table 18 of Appendix H.
@@ 4.1.4 Imputation
4.2 Classification and Anomaly detection
@@
Figure 3: Results of classification and anomaly detection. The results are averaged from several datasets. Higher accuracy and F1 score indicate better performance. $\ast$. indicates the Transformer-based models. See Table 19 and 20 in the Appendix H for full results.
@@ 4.3 Model Analysis
Figure 4: Visualization of representation on Time Image under Traffic dataset. More showcases in Figure 10,  11,  12.
@@ Appendix A Implementation Details
We evaluate the performance of different models for long-term forecasting on 8 well-established datasets, including Weather, Traffic, Electricity, Exchange, Solar-Energy, and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). Furthermore, we adopt PeMS and M4 datasets for short-term forecasting. We detail the descriptions of the dataset in Table 8.To comprehensively evaluate the model’s performance in time series analysis tasks, we further introduced datasets for classification and anomaly detection. The classification task is designed to test the model’s ability to capture coarse-grained patterns in time series data, while anomaly detection focuses on the recognition of fine-grained patterns. Specifically, we used 10 multivariate datasets from the UEA Time Series Classification Archive (2018) for the evaluation of classification tasks. For anomaly detection, we selected datasets such as SMD (2019), SWaT (2016), PSM (2021), MSL, and SMAP (2018). We detail the descriptions of the datasets for classification and anomaly detection in Table 9 and Table 10
@@
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2023), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022b), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023a), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2021) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2021), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@
All experiments were run three times, implemented in Pytorch (Paszke et al., 2019), and conducted on multi NVIDIA A100 80GB GPUs. We set the initial learning rate as $10^{-2}$ or $10^{-3}$ and used the ADAM optimizer (Kingma & Ba, 2015) with L2 loss for model optimization. And the batch size was set to be 512. By default, the TimeMixer++ contains 2 Mixing blocks. We set the number of resolutions $K$ to 3. Moreover, we set the number of MixerBlocks $L$ to 2. We choose the number of scales $M$ according to the time series length to balance performance and efficiency. To handle longer series in long-term forecasting, we set $M$ to 3. As for short-term forecasting with limited series length, we set $M$ to 1.
@@ Appendix B Details of Model design
As shown in Figure 9, at each scale $m$, the $K$ period-based representations, denoted as $\mathbf{z}_{m}^{(l,k)}$, are first weighted by their corresponding FFT amplitudes $\mathbf{A}_{f_{k}}$, which capture the importance of each period, and then summed to produce the final representation for that scale.
  • This process, repeated across all scales, yields a comprehensive embedding for each scale, capturing multi-resolution information.
@@ Appendix D Additional Representation Analysis
To evaluate the representational capabilities of TimeMixer++, we selected three datasets: Traffic, Electricity, and ETTm1, each exhibiting distinct periodic and trend characteristics. We conducted comprehensive analyses across various scales and resolutions. Initially, 1D convolution was employed to downsample the original time series, yielding different scales. Subsequently, frequency spectrum analysis was utilized to identify the primary frequency components within the time series, selecting the top three components by magnitude as the primary resolutions. This process transformed the original time series into a multi-scale time image. Time image decomposition was then applied to disentangle seasonality and trends from the original time image, resulting in distinct seasonality and trend images. Hierarchical mixing was performed to facilitate interactions across different scales. The visualization of the resulting representations is depicted in Figures 10, 11, and 12.

### v2

# Equal contribution    Project lead    Corresponding author

@@ 1 Introduction
At the core of TSPMs is their ability to recognize and generalize time series patterns inherent in time series data, enabling the model to uncover meaningful temporal structures and adapt to varying time series task scenarios. A line of research (Lai et al., 2018b; Zhao et al., 2017) has utilized recurrent neural networks (RNNs) to capture sequential patterns. However, these methods often struggle to capture long-term dependencies due to limitations like Markovian assumptions and inefficiencies. Temporal convolutional networks (TCNs) (Wu et al., 2023; Wang et al., 2023a; Liu et al., 2022a; Wang et al., 2024a) efficiently capture local patterns but face challenges with long-range dependencies (e.g., seasonality and trends) because of their fixed receptive fields. While some approaches reshape time series into 2D tensors based on frequency domain information (Wu et al., 2023) or downsample the time domain (Liu et al., 2022a), they fall short in comprehensively capturing long-range patterns. In contrast, transformer-based architectures (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2022b; Wang et al., 2022; Shi et al., 2024b) leverage token-wise self-attention to model long-range dependencies by allowing each token to attend to all others, overcoming the limitations of fixed receptive fields. However, unlike language tasks where tokens usually belong to distinct contexts, time series data often involve overlapping contexts at a single time point, such as daily, weekly, and seasonal patterns occurring simultaneously. This overlap makes it difficult to represent time series patterns effectively as tokens, posing challenges for transformer-based models in fully capturing the relevant temporal structures.
@@ 2 Related Work
Time Series Analysis. A pivotal aspect of time series analysis is the ability to extract diverse patterns from various time series while building powerful representations. This challenge has been explored across various model architectures. Traditional models like ARIMA (Anderson & Kendall, 1976) and STL (Cleveland et al., 1990) are effective for periodic and trend patterns but struggle with non-linear dynamics. Deep learning models, such as those by (Lai et al., 2018b) and (Zhao et al., 2017), capture sequential dependencies but face limitations with long-term dependencies. TCNs (Franceschi et al., 2019) improve local pattern extraction but are limited in capturing long-range dependencies. TimesNet (Wu et al., 2023) enhances long-range pattern extraction by treating time series as 2D signals, while MLP-based methods (Zeng et al., 2023; Ekambaram et al., 2023; Liu et al., 2023b; Wang et al., 2023c) offer simplicity and effectiveness. Transformer-based models like PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024) leverage self-attention to model long-range dependencies, demonstrating good forecasting performance. Given the strengths and limitations discussed above, there is a growing need for a TSPM capable of effectively extracting diverse patterns, adapting to various time series analytical tasks, and possessing strong generalization capabilities. As illustrated in Figure 1, TimeMixer++ meets this requirement by constructing robust representational capabilities, thereby demonstrating its potential for universal time series analysis.
@@
Hierarchical Time Series Modeling. Numerous methodologies have been advanced utilizing specialized deep learning architectures for time series analysis, with an emphasis on the decomposition and integration of temporal patterns. For example, several studies (Wu et al., 2021; Wang et al., 2024b; Zhou et al., 2022b; Luo et al., 2023; Wang et al., 2023b) utilize moving averages to discern seasonal and trend components, which are subsequently modeled using attention mechanisms (Wu et al., 2021; Zhou et al., 2022b; Shi et al., 2024a), convolutional networks (Wang et al., 2023a), or hierarchical MLP layers (Wang et al., 2024b; Wang, 2024). These components are individually processed prior to aggregation to yield the final output. Nonetheless, such approaches frequently depend on predefined and rigid operations for the disentanglement of seasonality and trends, thereby constraining their adaptability to complex and dynamic patterns. In contrast, as depicted in Figure 2, we propose a more flexible methodology that disentangles seasonality and trend directly within the latent space via dual-axis attention, thereby enhancing adaptability to a diverse range of time series patterns and task scenarios. Furthermore, by adopting a multi-scale, multi-resolution analytical framework (Mozer, 1991; Harti, 1993), we facilitate hierarchical interaction and integration across different scales and resolutions, substantially enhancing the effectiveness of time series modeling.
@@ 3.2 MixerBlock
where $\mathbf{A}=\{A_{f_{1}},\cdots,A_{f_{K}}\}$ represents the unnormalized amplitudes, $\{f_{1},\cdots,f_{K}\}$ are the top-$K$ frequencies, and $p_{k}=\left\lceil\frac{T}{f_{k}}\right\rceil,k\in\{1,\dots,K\}$ denotes the corresponding period lengths. Each time series representation $\mathbf{x}_{m}^{l}$ is then reshaped along the temporal dimension as follows:
@@
Time Image Decomposition. Time series patterns are inherently nested, with overlapping scales and periods. For example, weekly sales data reflects both daily shopping habits and broader seasonal trends. Conventional methods (Wu et al., 2021; Wang et al., 2024b) use moving averages across the entire series, often blurring distinct patterns. To address this, we utilize multi-resolution time images, where each image $\mathbf{z}_{m}^{(l,k)}\in\mathbb{R}^{p_{k}\times f_{\text{m,k}}\times d_{\text{model}}}$ encodes a specific scale and period, enabling finer disentanglement of seasonality and trend. By applying 2D convolution to these images, we capture long-range patterns and enhance temporal dependency extraction. Columns in each image correspond to time series segments within a period, while rows represent consistent time points across periods, facilitating dual-axis attention: column-axis attention ($\operatorname{Attention_{col}}$) captures seasonality within periods, and row-axis attention ($\operatorname{Attention_{row}}$) extracts trend across periods. Each axis-specific attention focuses on one axis, preserving efficiency by transposing the non-target axis to the batch dimension. For column-axis attention, queries, keys, and values $\mathbf{Q}_{\text{col}},\mathbf{K}_{\text{col}},\mathbf{V}_{\text{col}}\in\mathbb{R}^{f_{\text{m,k}}\times d_{\text{model}}}$ are computed via 2D convolution, which are shared across all images, and similarly for row-axis attention $\mathbf{Q}_{\text{row}},\mathbf{K}_{\text{row}},\mathbf{V}_{\text{row}}$. The seasonal and trend components are then computed as:
@@
Multi-scale Mixing. For each period $p_{k}$, we obtain $M+1$ seasonal time images and $M+1$ trend time images, denoted as $\{\mathbf{s}_{m}^{(l,k)}\}_{m=0}^{M}$ and $\{\mathbf{t}_{m}^{(l,k)}\}_{m=0}^{M}$, respectively. The 2D structure allows us to model both seasonal and trend patterns using 2D convolutional layers, which are more efficient and effective at capturing long-term dependencies than traditional linear layers (Wang et al., 2024b). For multi-scale seasonal time images, longer patterns can be interpreted as compositions of shorter ones (e.g., a yearly rainfall pattern formed by monthly changes). Therefore, we mix the seasonal patterns from fine-scale to coarse-scale. To facilitate this bottom-up information flow, we apply the 2D convolutional layers at the $m$-th scale in a residual manner, formalized as:
@@ 4.1.1 Long-term forecasting
Setups. Long-term forecasting is pivotal for strategic planning in areas such as weather prediction, traffic management, and energy utilization. To comprehensively assess our model’s effectiveness over extended periods, we perform experiments on 8 widely-used real-world datasets, including the four subsets of the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), as well as Weather, Solar-Energy, Electricity, and Traffic, consistent with prior benchmarks set by Zhou et al. (2021b); Liu et al. (2022a).
Table 1: Long-term forecasting results. We average the results across 4 prediction lengths: $\{96,192,336,720\}$. The best performance is highlighted in red, and the second-best is underlined. Full results can be found in Appendix H.
@@ 4.1.2 Univariate short-term forecasting
Table 2: Univariate short-term forecasting results, averaged across all M4 subsets. Full results are available in Appendix H
@@
Results. Table 2 demonstrates that TimeMixer++ significantly outperforms state-of-the-art models across all metrics. Compared to iTransformer, it reduces SMAPE by 9.7% and MASE by 15.7%, with even larger improvements over TiDE, achieving up to a 23.3% reduction in MASE. Additionally, TimeMixer++ records the lowest OWA, outperforming TimesNet by 3.5% and iTransformer by 11.6%.
@@ 4.1.3 Multivariate short-term forecasting
Table 3: Results of multivariate short-term forecasting, averaged across all PEMS datasets. Full results can be found in Table 18 of Appendix H.
@@ 4.1.4 Imputation
4.1.7 Classification and Anomaly detection
@@
Figure 3: Results of classification and anomaly detection. The results are averaged from several datasets. Higher accuracy and F1 score indicate better performance. $\ast$. indicates the Transformer-based models. See Table 21 and 22 in the Appendix H for full results.
@@ 4.2 Model Analysis
Figure 4: Visualization of representation on Time Image under Traffic dataset. More showcases in Figure 10,  11,  13.
@@ 5 Conclusion
Acknowledgement
M. Jin was supported in part by the NVIDIA Academic Grant Program and CSIRO – National Science Foundation (US) AI Research Collaboration Program.
@@ Appendix A Implementation Details
We evaluate the performance of different models for long-term forecasting on 8 well-established datasets, including Weather, Traffic, Electricity, Exchange, Solar-Energy, and ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). Furthermore, we adopt PeMS and M4 datasets for short-term forecasting. We detail the descriptions of the dataset in Table 8.To comprehensively evaluate the model’s performance in time series analysis tasks, we further introduced datasets for classification and anomaly detection. The classification task is designed to test the model’s ability to capture coarse-grained patterns in time series data, while anomaly detection focuses on the recognition of fine-grained patterns. Specifically, we used 10 multivariate datasets from the UEA Time Series Classification Archive (2018) for the evaluation of classification tasks. For anomaly detection, we selected datasets such as SMD (2019), SWaT (2016), PSM (2021), MSL, and SMAP (2018). We detail the descriptions of the datasets for classification and anomaly detection in Table 9 and Table 10.
@@
To assess the effectiveness of our method across various tasks, we select 27 advanced baseline models spanning a wide range of architectures. Specifically, we utilize CNN-based models: MICN (2023a), SCINet (2022a), and TimesNet (2023); MLP-based models: TimeMixer (2024b), LightTS (2022a), and DLinear (2023); RMLP&RLinear (2023a) and Transformer-based models: iTransformer (2023a), PatchTST (2023), Crossformer (2023), FEDformer (2022b), Stationary (2022b), Autoformer (2021), and Informer (2021b). These models have demonstrated superior capabilities in temporal modeling and provide a robust framework for comparative analysis. For specific tasks, TiDE (2023a), FiLM (2022a), N-HiTS (2023), and N-BEATS (2019) address long- or short-term forecasting; Anomaly Transformer (2021) and MTS-Mixers (2023b) target anomaly detection; while Rocket (2023a), LSTNet (2018c), LSSL (2021), and Flowformer (2022) are utilized for classification. Few/zero-shot forecasting tasks employ ETSformer (2022), Reformer (2020), and LLMTime (2023).
@@
All experiments were run three times, implemented in Pytorch (Paszke et al., 2019), and conducted on multi NVIDIA A100 80GB GPUs. We set the initial learning rate as a range from $10^{-3}$ to $10^{-1}$ and used the ADAM optimizer (Kingma & Ba, 2015) with L2 loss for model optimization. And the batch size was set to be 512. We set the number of resolutions $K$ to range from 1 to 5. Moreover, we set the number of MixerBlocks $L$ to range from 1 to 3. We choose the number of scales $M$ according to the time series length to balance performance and efficiency. To handle longer series in long-term forecasting, we usually set $M$ to 3. As for short-term forecasting with limited series length, we usually set $M$ to 1. We pretrained the model with learning rate decay after linear warm-up. For baselines under the same experimental settings as our main study, we directly report the results from TimesNet (Wu et al., 2023). In scenarios where experimental settings differed or tasks were not previously implemented, we reproduced the baseline results referring to the benchmark framework from the Time-Series Library  [^2] https://github.com/thuml/Time-Series-Library. The source code and pretrained model will be provided in GitHub (https://github.com/kwuking/TimeMixer).
@@ Appendix B Details of Model design
As shown in Figure 9, at each scale $m$, the $K$ period-based representations, denoted as $\mathbf{z}_{m}^{(l,k)}$, are first weighted by their corresponding FFT amplitudes $\mathbf{A}_{f_{k}}$, which capture the importance of each period, and then summed to produce the final representation for that scale. This process, repeated across all scales, yields a comprehensive embedding for each scale, capturing multi-resolution information.
@@ Appendix D Additional Representation Analysis
To evaluate the representational capabilities of TimeMixer++, we selected three datasets: Traffic, Electricity, and ETTm1, each exhibiting distinct periodic and trend characteristics. We conducted comprehensive analyses across various scales and resolutions. Initially, 1D convolution was employed to downsample the original time series, yielding different scales. Subsequently, frequency spectrum analysis was utilized to identify the primary frequency components within the time series, selecting the top three components by magnitude as the primary resolutions. This process transformed the original time series into a multi-scale time image. Time image decomposition was then applied to disentangle seasonality and trends from the original time image, resulting in distinct seasonality and trend images. Hierarchical mixing was performed to facilitate interactions across different scales. The visualization of the resulting representations is depicted in Figures 10, 11, and 13.
@@
Figure 12: Representation analysis in four tasks. For each model, the centered kernel alignment (CKA) similarity is computed between representations from the first and the last layers.
@@
As shown in Figure 12, TimeMixer++ consistently excels across all four tasks, achieving state-of-the-art performance in each scenario. The model effectively adapts its representation transformations to meet the demands of different tasks. Whether preserving representations or enabling significant changes across layers, TimeMixer++ demonstrates its versatility and robustness in handling a wide range of time-series analysis challenges.